# Cycles in spherical Deligne complexes and application to $K(\pi,1)$-conjecture for Artin groups

### v1

@@ 11 $K(\pi,1)$ for some higher-dimensional families
Suppose $\Lambda$ is a Dynkin diagram of type $B_{n},\widetilde{C}_{n},H_{3}$ or $F_{4}$. Let $\Lambda^{\prime}$ be a linear subgraph with three vertices. Then $\Delta_{\Lambda,\Lambda^{\prime}}$ is weakly flag.
@@
First assume $\Lambda$ is of type $B_{n}$. We label consecutive vertices of $\Lambda$ by $\{s_{i}\}_{i=1}^{n}$ with $m_{s_{n-1},s_{n}}=4$. Then vertices of $\Delta_{\Lambda}$ with the order induced from $s_{1}<s_{2}<\cdots<s_{n}$ is upward flag. Suppose $\Lambda^{\prime}$ has vertices $\{s_{i},s_{i+1},s_{i+2}\}$. Then we need to show:
@@
Suppose $\Lambda$ is a Dynkin diagram belonging to one of the following seven families in Figure 16. Then $A_{\Lambda}$ satisfies the $K(\pi,1)$-conjecture.
Figure 16: Seven families.
@@
Let $\mathcal{C}_{2},\mathcal{C}_{3},\mathcal{C}_{4}$ be the families in first row right, second row left, and second two right respectively in Figure 16. All families in Figure 16 except $\mathcal{C}_{2},\mathcal{C}_{3}$ and $\mathcal{C}_{4}$ follow from Proposition 11.1 (the choice of $\Lambda^{\prime}$ are indicated in the thickened subgraphs). Let $\mathcal{C}_{2}$ be the remaining family of Dynkin diagrams. We will verify the assumptions of [Hua23, Corollary 7.3] with $\mathcal{C}_{1}=\{\widetilde{B}_{3}\}$. Assumptions $2,3$ of [Hua23, Corollary 7.3] are clear. Take $\Lambda\in\mathcal{C}_{2}$ and let $\Lambda^{\prime}$ be the thickened subgraph of $\Lambda$. Then $\Lambda\setminus\{s\}$ is spherical for $s=b_{1}$ or $a$, and $\Lambda\setminus\{s\}$ belongs to the family on the left side of second row for $s=b_{2}$ or $b_{3}$. Thus Assumption $4$ of [Hua23, Corollary 7.3] follows. It remains to show $\Delta_{\Lambda,\Lambda^{\prime}}$ is contractible. For $i=1,2$, let $\Lambda_{i}=\Lambda\setminus\{b_{i}\}$. Let $\Lambda^{\prime}_{i}$ be the subgraph spanned by $\Lambda_{i}\cap\Lambda^{\prime}$ and $\{b_{4}\}$. Then by the same argument as in the proof of Proposition 11.1, we know $\Delta_{\Lambda_{i},\Lambda^{\prime}_{i}}$ satisfies the assumptions of Proposition 3.35. Thus Proposition 8.1 (1) implies that $\Delta_{\Lambda_{i},\Lambda_{i}\cap\Lambda^{\prime}}$ is a bowtie free, upward flag poset with its vertex set endowed with the order induced from $b_{i}<a<b_{1}$. Thus Assumption 1 of Proposition 3.35 holds for $\Delta_{\Lambda,\Lambda^{\prime}}$. Assumption 2 also holds, by Theorem 3.16. Thus $\Delta_{\Lambda,\Lambda^{\prime}}$ is contractible by Proposition 3.35.
Treating families $\mathcal{C}_{3}$ and $\mathcal{C}_{4}$ reduces to showing $\Delta_{\Lambda,\Lambda^{\prime}}$ is contractible, where $\Lambda$ is a diagram in $\mathcal{C}_{3}$ or $\mathcal{C}_{4}$, and $\Lambda^{\prime}$ is the thickened $\widetilde{A}_{3}$ subdiagram in Figure 16. For a vertex $s\in\Lambda^{\prime}$, let $\Lambda_{s}$ be the component of $\Lambda\setminus\{s\}$ containing $\Lambda^{\prime}\setminus\{s\}$. By Theorem 3.27 and Lemma 3.4, it suffices to show $\Delta_{\Lambda_{s},\lambda_{s}\cap\Lambda^{\prime}}$ is bowtie free for each $s\in\Lambda^{\prime}$. Note that $\Lambda_{s}$ is of type $B_{n},D_{n},\widetilde{B}_{4}$ or $\widetilde{D}_{4}$. The $B_{n}$ and $D_{n}$ case follows from Theorem 3.17. If $\Lambda_{s}$ is of type $\widetilde{B}_{4}$, then by Theorem 10.1 and Theorem 3.22, $\Delta_{\Lambda_{s}}$ satisfies the assumptions of Proposition 3.32. By Proposition 8.1 (3) (the “in addition” assumption follows from Theorem 10.1) and Lemma 3.20, $\Delta_{\Lambda_{s},\Lambda_{s}\cap\Lambda^{\prime}}$ is bowtie free. If $\Lambda_{3}$ is of type $\widetilde{D}_{4}$, then by Theorem 10.1, $\Delta_{\Lambda_{s}}$ satisfies the assumptions of Proposition 3.34. Thus by Proposition 8.8 $\Delta_{\Lambda_{s},\Lambda_{s}\cap\Lambda^{\prime}}$ is bowtie free. ∎
@@
Let $W_{\Lambda}$ be a reflection group acting on $\mathbb{H}^{n}$ with $n\leq 4$ such that the fundamental domain is a finite volume non-compact simplex, and $\Lambda$ is the Dynkin diagram. Then the $K(\pi,1)$-conjecture holds for $A_{\Lambda}$.
Figure 17: Dynkin diagrams for 4-dimensional quasi-Lanner groups.
@@
The $n\leq 2$ case follows from [CD95a]. The $n=3$ case follows from Corollary 9.3. For $n=4$, the only possible Dynkin diagrams are shown in Figure 17. The $K(\pi,1)$-conjecture for diagrams (1), (2) and (4) are proved in [Hua23]. The remaining diagrams are consequences of Proposition 11.1 (the choice of $\Lambda^{\prime}$ is the thickened subgraphs in Figure 17). ∎
Theorem 11.5.
Suppose $\Lambda$ is a tree Dynkin diagram with a vertex $s\in\Lambda$ such that
  1. 1.
    each edge containing $s$ has label $=3$;
  2. 2.
    each component of $\Lambda\setminus\{s\}$ either contains only one vertex, or contains only one edge labeled by $4$.
Then $A_{\Lambda}$ satisfies the $K(\pi,1)$-conjecture.

### v3

@@ 12 $K(\pi,1)$ for some higher-dimensional families
Suppose $\Lambda$ is a Coxeter diagram of type $B_{n},\widetilde{C}_{n},H_{3}$ or $F_{4}$. Let $\Lambda^{\prime}$ be a linear subgraph with three nodes. Then $\Delta_{\Lambda,\Lambda^{\prime}}$ is weakly flag.
@@
First assume $\Lambda$ is of type $B_{n}$. We label consecutive nodes of $\Lambda$ by $\{s_{i}\}_{i=1}^{n}$ with $m_{s_{n-1},s_{n}}=4$. Then vertices of $\Delta_{\Lambda}$ with the order induced from $s_{1}<s_{2}<\cdots<s_{n}$ is upward flag. Suppose $\Lambda^{\prime}$ has nodes $\{s_{i},s_{i+1},s_{i+2}\}$. Then we need to show:
@@
The downward flag part of Conjecture 2.34 holds true when $n=3$, and $A_{S}$ is of type $A_{n},B_{n},H_{3},F_{4}$.
@@
This is a combination of Lemma 8.2, Lemma 12.2, Theorem 5.6, Theorem 7.1 and Proposition 3.2. ∎
@@
Suppose $\Lambda$ is a Coxeter diagram belonging to one of the following families in Figure 15. Then $A_{\Lambda}$ satisfies the $K(\pi,1)$-conjecture.
Figure 15: Several families.
@@
Let $\mathcal{C}_{2},\mathcal{C}_{3},\mathcal{C}_{4}$ be the families in first row right, second row left, and second two right respectively in Figure 15. All families in Figure 15 except $\mathcal{C}_{2},\mathcal{C}_{3}$ and $\mathcal{C}_{4}$ follow from Proposition 12.1 (the choice of $\Lambda^{\prime}$ are indicated in the thickened subgraphs). Let $\mathcal{C}_{2}$ be the remaining family of Coxeter diagrams. We will verify the assumptions of [Hua24a, Corollary 7.3] with $\mathcal{C}_{1}=\{\widetilde{B}_{3}\}$. Assumptions $2,3$ of [Hua24a, Corollary 7.3] are clear. Take $\Lambda\in\mathcal{C}_{2}$ and let $\Lambda^{\prime}$ be the thickened subgraph of $\Lambda$. Then $\Lambda\setminus\{s\}$ is spherical for $s=b_{1}$ or $a$, and $\Lambda\setminus\{s\}$ belongs to the family on the left side of second row for $s=b_{2}$ or $b_{3}$. Thus Assumption $4$ of [Hua24a, Corollary 7.3] follows. It remains to show $\Delta_{\Lambda,\Lambda^{\prime}}$ is contractible. For $i=1,2$, let $\Lambda_{i}=\Lambda\setminus\{b_{i}\}$. Let $\Lambda^{\prime}_{i}$ be the subgraph spanned by $\Lambda_{i}\cap\Lambda^{\prime}$ and $\{b_{4}\}$. Then by the same argument as in the proof of Proposition 12.1, we know $\Delta_{\Lambda_{i},\Lambda^{\prime}_{i}}$ satisfies the assumptions of Proposition 8.3. Thus Proposition 9.1 (1) implies that $\Delta_{\Lambda_{i},\Lambda_{i}\cap\Lambda^{\prime}}$ is a bowtie free, upward flag poset with its vertex set endowed with the order induced from $b_{i}<a<b_{1}$. Thus Assumption 1 of Proposition 8.3 holds for $\Delta_{\Lambda,\Lambda^{\prime}}$. Assumption 2 also holds, by Theorem 2.38. Thus $\Delta_{\Lambda,\Lambda^{\prime}}$ is contractible by Proposition 8.3.
Treating families $\mathcal{C}_{3}$ and $\mathcal{C}_{4}$ reduces to showing $\Delta_{\Lambda,\Lambda^{\prime}}$ is contractible, where $\Lambda$ is a diagram in $\mathcal{C}_{3}$ or $\mathcal{C}_{4}$, and $\Lambda^{\prime}$ is the thickened $\widetilde{A}_{3}$ subdiagram in Figure 15. For a node $s\in\Lambda^{\prime}$, let $\Lambda_{s}$ be the component of $\Lambda\setminus\{s\}$ containing $\Lambda^{\prime}\setminus\{s\}$. By Theorem 2.18 and Lemma 2.23, it suffices to show $\Delta_{\Lambda_{s},\lambda_{s}\cap\Lambda^{\prime}}$ is bowtie free for each $s\in\Lambda^{\prime}$. Note that $\Lambda_{s}$ is of type $B_{n},D_{n},\widetilde{B}_{4}$ or $\widetilde{D}_{4}$. The $B_{n}$ and $D_{n}$ case follows from Theorem 2.35. If $\Lambda_{s}$ is of type $\widetilde{B}_{4}$, then by Theorem 11.3 and Theorem 2.37, $\Delta_{\Lambda_{s}}$ satisfies the assumptions of Proposition 2.24. By Proposition 9.1 (3) (the “in addition” assumption follows from Theorem 11.3) and Lemma 2.30, $\Delta_{\Lambda_{s},\Lambda_{s}\cap\Lambda^{\prime}}$ is bowtie free. If $\Lambda_{3}$ is of type $\widetilde{D}_{4}$, then by Theorem 11.3, $\Delta_{\Lambda_{s}}$ satisfies the assumptions of Proposition 2.26. Thus by Proposition 9.2 $\Delta_{\Lambda_{s},\Lambda_{s}\cap\Lambda^{\prime}}$ is bowtie free. ∎
Corollary 12.5.
Let $W_{\Lambda}$ be a reflection group acting on $\mathbb{H}^{n}$ with $n\leq 4$ such that the fundamental domain is a finite volume non-compact simplex, and $\Lambda$ is the Coxeter diagram. Then the $K(\pi,1)$-conjecture holds for $A_{\Lambda}$.
Figure 16: Coxeter diagrams for 4-dimensional quasi-Lanner groups.